OpenStudy (anonymous):

independent events a&b, p(A) = 0.50 , p(B) = 0.45 find P(A or B)

OpenStudy (anonymous):

0.95?

OpenStudy (anonymous):

P(a or b) p(a) + p(b) = 0.5 + 0.45 = 0.95